a)Define the following pertaining to inventory systems:
i.Deterministic Inventory Model
ii.Economic Order Quantity
iii.Reorder Point
iv.Lead Time
v.Backorder

b)Tele-Reco is a new specialty store that sells mobile phones, laptops and other computer accessories. A new Japanese-manufactured mobile phone reorder costs Tele-Reco $600 per unit. Tele-Reco’s annual holding cost rate is 22%. Ordering costs are estimated to be $70 per order.
i.If the demand for the new mobile phone is expected to be constant with a rate of 20 units per month, what is the recommended annual order quantity for the new mobile phone?
ii.What is the estimated annual inventory cost associated with this product?
iii.How many orders will be placed per year?
iv.With 250 working days, what is the cycle time for this product?
1
Expert's answer
2020-06-15T18:17:15-0400

a. i.Deterministic models of inventory control are used to determine the optimal inventory of a single item when demand is mostly largely obscure

ii. Economic order quantity (EOQ) is the ideal order quantity a company should purchase to minimize inventory costs such as holding costs, shortage costs, and order costs. 

iii.The reorder point (ROP) is the level of inventory which triggers an action to replenish that particular inventory stock

iv.The lead time is the delay applicable for inventory control purposes. 

v. Backorder means to place an order for a product that is temporarily out of stock.

b)i.Annual order quantity="\sqrt{(2 \times \, Co\, \times \,D)/C}"

Where:

  • Co is the cost of placing one order
  • D is the annual demand
  • C is the annual cost of holding one unit of inventory
  • ="\sqrt{(2 \cdot 70)/0.22}=25.23"

ii.ii.annual holding cost=ordering cost

=70

iii.240/39=6 orders per year

iv,250/6=42